# LinguDistill : Recovering Linguistic Ability in Vision-Language Models via Selective Cross-Modal Distillation

### 01 Apr 2026 v1

@@ Abstract
Adapting pretrained language models (LMs) into vision-language models (VLMs) can degrade their native linguistic capability due to representation shift and cross-modal interference introduced during multimodal adaptation. Such loss is difficult to recover, even with targeted task-specific fine-tuning using standard objectives. Prior recovery approaches typically introduce additional modules that act as intermediate alignment layers to maintain or isolate modality-specific subspaces, which increases architectural complexity, adds parameters at inference time, and limits flexibility across models and settings. We propose LinguDistill , an adapter-free distillation method that restores linguistic capability by utilizing the original frozen LM as a teacher. We overcome the key challenge of enabling vision-conditioned teacher supervision by introducing layer-wise KV-cache sharing, which exposes the teacher to the student’s multimodal representations without modifying the architecture of either model. We then selectively distill the teacher’s strong linguistic signal on language-intensive data to recover language capability, while preserving the student’s visual grounding on multimodal tasks. As a result, LinguDistill recovers $\sim$10% of the performance lost on language and knowledge benchmarks, while maintaining comparable performance on vision-heavy tasks. Our findings demonstrate that linguistic capability can be recovered without additional modules, providing an efficient and practical solution to modality-specific degradation in multimodal models.
@@ 5 Results
Table 3: Benchmark results across all variants. Color encodes direction relative to the step-0 baseline: >>10% regression, 5–10% regression, 1–5% regression, $\pm$1% comparable, 2–3% gain, 3–5% gain, >>5% gain. It is observable that we are able to recover language-heavy capability the best while having minimum regression on visual-heavy tasks, compared to non-selective distillation and standard fine-tuning.
@@
Language and knowledge-heavy
AI2D 0.440 0.416 0.413 0.490 0.502 0.507
COCO 2017 Captioning 0.800 0.673 0.742 0.843 0.787 0.866
ScienceQA 0.590 0.592 0.581 0.650 0.679 0.676
ARC Easy 0.605 0.540 0.542 0.548 0.601 0.621
ARC Challenge 0.322 0.279 0.281 0.270 0.296 0.318
HellaSwag 0.405 0.326 0.329 0.345 0.376 0.394
Document and OCR
DocVQA 0.769 0.767 0.766 0.640 0.709 0.740
InfographicVQA 0.357 0.282 0.272 0.290 0.318 0.330
OCRBench 0.760 0.726 0.722 0.452 0.510 0.600
@@
RealWorldQA 0.523 0.500 0.488 0.450 0.494 0.490
MME Cognition 302 229 261 241 308 300
MME Perception 1259 1027 1070 1046 1109 1173
MMMU 0.320 0.310 0.302 0.283 0.288 0.283
MMStar 0.360 0.359 0.358 0.350 0.355 0.349
@@
nanoVLM-lang restricts training to the 8 language-heavy sources, removing all OCR and document data. This partially reduces the COCO regression ($7.3\%$ vs. $15.9\%$) but fails to recover text-only benchmarks. HellaSwag still drops to 0.329 ($18.8\%$), ARC Easy to 0.542 ($10.4\%$), and MME cognition to 261 ($13.6\%$). This gives conclusion that Data filtering alone cannot correct the backbone drift caused by multimodal fine-tuning.
@@
Uniform distillation partially reduces language regression but damages vision tasks.
@@
distill-full applies the KD signal uniformly across all sources with $\alpha{=}0.5$ and T=2T{=}2. Language regression is partially mitigated: COCO rises to 0.843 ($5.4\%$ over nanoVLM), AI2D to 0.490 ($11.4\%$ over vanilla), and ScienceQA to 0.650 ($10.2\%$ over vanilla). However, OCR and document tasks are severely damaged: OCRBench drops from 0.760 to 0.452 ($40.5\%$) and DocVQA from 0.769 to 0.640 ($16.8\%$). This signals that supervision from pure teacher LM with limited visual grounding capability, even with KV-cache sharing, actively mislead the student on pixel-level tasks. Text-only reasoning benchmarks also remain degraded: HellaSwag drops to 0.345 ($14.8\%$), ARC Easy to 0.548 ($9.4\%$), and ARC Challenge to 0.270 ($16.1\%$). We conclude that uniform distillation mimics pure teacher forcing behavior, suppressing the student’s own supervisory signal entirely and leaving no regularization to retain original nanoVLM’s capability, which leads to the mentioned task-specific catastrophic forgetting.
@@
Selective distillation achieves the best language recovery with contained vision degradation.
@@
LinguDistill trains on all 17 sources (full subset) but applies $\alpha{=}0$ to the 9 OCR/document sources, selectively absorbing KD signal only to language-heavy examples. This results in the strongest language gains across all variants: ScienceQA 0.676 ($14.6\%$ over vanilla), AI2D 0.507 ($15.2\%$), COCO 0.866 ($8.3\%$), and MME cognition 300 ($0.7\%$, essentially unchanged from vanilla). Text-only reasoning benchmarks also show substantially reduced regression: HellaSwag drops to only 0.394 ($2.7\%$), ARC Easy to 0.621 ($2.6\%$ over vanilla), and ARC Challenge to 0.318 ($1.2\%$). These benchmarks have no visual component and directly reflect linguistic capability. Vision task degradation is substantially reduced relative to uniform distillation and standard finetuning: DocVQA drops $3.8\%$ (vs. $16.8\%$ under distill-full), InfoVQA $7.6\%$ (vs. $18.8\%$), and OCRBench $21.1\%$ (vs. $40.5\%$). We further analyze the drop in OCRBench performance and find that these tasks rely on the student’s visual grounding, which the text-only teacher cannot provide (Appendix E). MMMU regresses by $11.6\%$ across all distillation variants; we attribute this to our single-image, 1024-token training constraint rather than the distillation objective itself (Appendix D).

### 25 Apr 2026 v3

@@ Abstract
Adapting pretrained language models (LMs) into vision-language models (VLMs) can degrade their native linguistic capability due to representation shift and cross-modal interference introduced during multimodal adaptation. Such loss is difficult to recover, even with targeted task-specific fine-tuning using standard objectives. Prior recovery approaches typically introduce additional modules that act as intermediate alignment layers to maintain or isolate modality-specific subspaces, which increases architectural complexity, adds parameters at inference time, and limits flexibility across models and settings. We propose LinguDistill , an adapter-free distillation method that restores linguistic capability by utilizing the original frozen LM as a teacher. We overcome the key challenge of enabling vision-conditioned teacher supervision by introducing layer-wise KV-cache sharing, which exposes the teacher to the student’s multimodal representations without modifying the architecture of either model. We then selectively distill the teacher’s strong linguistic signal on language-intensive data to recover language capability, while preserving the student’s visual grounding on multimodal tasks. As a result, LinguDistill recovers $\sim$10% of the performance lost on language and knowledge benchmarks, while maintaining comparable performance on vision-heavy tasks. Subject to further scaling tests, our findings suggest that linguistic capability of multimodal models can be recovered without extra modules, providing an efficient solution to modality degradation.
@@ 5 Results
Table 3: Benchmark results across variants. Values are absolute scores; colored percentages denote change vs. nanoVLM-full. Average rows are manually color-scaled for perceptual consistency: small changes are subtle, larger changes progressively emphasized. Selective KD (LinguDistill) achieves the strongest gains coupled with a controlled perception ability degradation.
@@
Language-heavy
AI2D 0.440 0.416 0.413 (-0.7%) 0.490 (+17.8%) 0.502 (+20.7%) 0.507 (+21.9%)
COCO 0.800 0.673 0.742 (+10.3%) 0.843 (+25.3%) 0.787 (+16.9%) 0.866 (+28.7%)
ScienceQA 0.590 0.592 0.581 (-1.9%) 0.650 (+9.8%) 0.679 (+14.7%) 0.676 (+14.2%)
ARC Easy 0.605 0.540 0.542 (+0.4%) 0.548 (+1.5%) 0.601 (+11.3%) 0.621 (+15.0%)
ARC Challenge 0.322 0.279 0.281 (+0.7%) 0.270 (-3.2%) 0.296 (+6.1%) 0.318 (+14.0%)
HellaSwag 0.405 0.326 0.329 (+0.9%) 0.345 (+5.8%) 0.376 (+15.3%) 0.394 (+20.9%)
Average 0.527 0.471 0.481(+2.1%) 0.524(+11.3%) 0.540(+14.6%) 0.564(+19.7%)
Document & OCR
DocVQA 0.769 0.767 0.766 (-0.1%) 0.640 (-16.6%) 0.709 (-7.6%) 0.740 (-3.5%)
InfographicVQA 0.357 0.282 0.272 (-3.5%) 0.290 (+2.8%) 0.318 (+12.8%) 0.330 (+17.0%)
OCRBench 0.760 0.726 0.722 (-0.6%) 0.452 (-37.7%) 0.510 (-29.7%) 0.600 (-17.4%)
Average 0.629 0.592 0.587(-0.8%) 0.461(-22.1%) 0.512(-13.5%) 0.557(-5.9%)
@@
MME Cognition 302 229 261 (+14.0%) 241 (+5.2%) 308 (+34.5%) 300 (+31.0%)
MME Perception 1259 1027 1070 (+4.2%) 1046 (+1.8%) 1109 (+8.0%) 1173 (+14.2%)
MME Avg 780.5 628.0 665.5(+6.0%) 643.5(+2.5%) 708.5(+12.8%) 736.5(+17.3%)
RealWorldQA 0.523 0.500 0.488 (-2.4%) 0.450 (-10.0%) 0.494 (-1.2%) 0.490 (-2.0%)
MMMU 0.320 0.310 0.302 (-2.6%) 0.283 (-8.7%) 0.288 (-7.1%) 0.283 (-8.7%)
MMStar 0.360 0.359 0.358 (-0.3%) 0.350 (-2.5%) 0.355 (-1.1%) 0.349 (-2.8%)
Non-MME Avg 0.401 0.390 0.383(-1.8%) 0.361(-7.4%) 0.379(-2.8%) 0.374(-4.1%)
@@
nanoVLM-lang restricts training to the 8 language-heavy sources, removing all OCR and document data. This partially improves COCO ($10.3\%$ relative to nanoVLM-full) but shows inconsistent gains on reasoning benchmarks. HellaSwag improves only marginally to 0.329 ($0.9\%$), ARC Easy to 0.542 ($0.4\%$), while ARC Challenge remains limited at 0.281 ($0.7\%$). MME cognition improves to 261 ($14.0\%$), indicating partial recovery. However, we also observe slight regressions on AI2D ($0.7\%$) and ScienceQA ($1.9\%$), which are themselves language- and knowledge-intensive benchmarks. These degradations are counterintuitive, as such tasks should benefit from restricting training to language-heavy data, and suggest that the underlying representation drift is not resolved by data filtering alone. This suggests that while filtering helps reduce some degradation, it is insufficient to fully correct the backbone drift introduced by multimodal fine-tuning.
@@
Uniform distillation partially improves language performance but damages vision tasks.
@@
distill-full applies the KD signal uniformly across all sources with $\alpha{=}0.5$ and T=2T{=}2. Language benchmarks improve substantially over the fine-tuning baseline: COCO rises to 0.843 ($25.3\%$), AI2D to 0.490 ($17.8\%$), and ScienceQA to 0.650 ($9.8\%$). Text-only reasoning benchmarks also improve modestly: HellaSwag reaches 0.345 ($5.8\%$), ARC Easy 0.548 ($1.5\%$), while ARC Challenge remains slightly degraded at 0.270 ($3.2\%$). However, OCR and document tasks are severely damaged: OCRBench drops from 0.726 to 0.452 ($37.7\%$) and DocVQA from 0.767 to 0.640 ($16.6\%$). This signals that supervision from a pure language teacher, even with KV-cache sharing, misguides the student on pixel-level tasks lacking visual grounding. We conclude that uniform distillation over-regularizes the model toward language behavior, suppressing modality-specific signals and leading to task-specific degradation.
@@
Selective distillation achieves the best language gains with contained vision degradation.
@@
LinguDistill trains on all 17 sources (full subset) but applies $\alpha{=}0$ to the 9 OCR/document sources, selectively absorbing KD signal only for language-heavy examples. This results in the strongest improvements across all language benchmarks: ScienceQA reaches 0.676 ($14.2\%$), AI2D 0.507 ($21.9\%$), COCO 0.866 ($28.7\%$), and MME cognition 300 ($31.0\%$). Text-only reasoning benchmarks also show consistent gains: HellaSwag improves to 0.394 ($20.9\%$), ARC Easy to 0.621 ($15.0\%$), and ARC Challenge to 0.318 ($14.0\%$). These benchmarks have no visual component and directly reflect linguistic capability.
Vision task degradation is substantially reduced compared to uniform distillation: DocVQA drops only $3.5\%$ (vs. $16.6\%$), while OCRBench drops $17.4\%$ (vs. $37.7\%$). InfographicVQA instead improves to 0.330 ($17.0\%$), indicating better balance between modalities. We further analyze OCRBench degradation and find that these tasks rely on visual grounding signals unavailable to the text-only teacher (Appendix E). MMMU remains degraded at $8.7\%$ across distillation variants; we attribute this to the single-image, 1024-token training constraint rather than the distillation objective itself (Appendix D).
Overall, selective distillation achieves a superior trade-off: it maximizes language gains while preserving visual capability significantly better than uniform distillation.